OpenStudy (anonymous):

MEDALS!!! The length of the hypotenuse of a right triangle is 16 inches. If the length of one leg is 5 inches, what is the approximate length of the other leg? 10.5 inches 11.0 inches 15.2 inches 16.8 inches

OpenStudy (tkhunny):

Have you considered he Pythagorean Theorem?

OpenStudy (ffrocks3):

15.2 Inches

OpenStudy (ffrocks3):

Do You have anymore questions? Just ask here and I can try to answer for you :)

OpenStudy (anonymous):

@Ffrocks3 thank man! You are such a big help!

OpenStudy (ffrocks3):

Np! ;)

OpenStudy (anonymous):

@tkhunny im doing it now!

OpenStudy (ffrocks3):

Want me to tell you how?

OpenStudy (anonymous):

@tkhunny i mean im in that lesson.

OpenStudy (ffrocks3):

Using the formula c=a2+b2 Solving forb b=c2﹣a2=162﹣52≈15.19868

OpenStudy (anonymous):

@Ffrocks3 yes!

OpenStudy (ffrocks3):

\[C=\sqrt{a ^{2}+b ^{2}}\] Solving for b \[b = \sqrt{c ^{2}-a ^{2}} = \sqrt{16^{2}-5^{2}} \]

OpenStudy (tkhunny):

@Ffrocks3 Let's get a little better at notation, please. c^2 = a^2 + b^2 is much more clear. Also, NEVER as if the OP want's to see how. Always get the OP to do the work.

OpenStudy (ffrocks3):

Yea sorry about that I didn't know I will now :)

OpenStudy (tkhunny):

Also, don't change notation in the middle of the problem. "C" and "c" are not the same thing.

OpenStudy (ffrocks3):

yep

OpenStudy (anonymous):

@Ffrocks3 thank you for helping me understand this problem. @tkhunny you too.